OpenStudy (anonymous):

please help! i'll give you a cookie c; What is the value of the expression 7^-3 A. -21 B. -1/343 C. 1/343 D. 343

OpenStudy (whpalmer4):

what is \(7^{-3}=\) property of exponents you need to remember: \[x^{-a} = \frac{1}{x^{a}}\] and \[\frac{1}{x^{-a}} = x^a\]

OpenStudy (whpalmer4):

I like ginger spice cookies, please :-)

OpenStudy (whpalmer4):

What did you get for the answer?

OpenStudy (anonymous):

from what i got, i think the answer is B...?

OpenStudy (whpalmer4):

oh, I'm sorry, I'm not going to be able to share my cookies with you....

OpenStudy (whpalmer4):

compare your problem with my formula: \[7^{-3}\]\[x^{-a}\] what is the value of \(x\)? What is the value of \(a\)?

OpenStudy (anonymous):

uh.. im not sure.... all i did was -7*-7*-7 i dont even know if thats what i was supposed to do but i tried it. (im new to this). but i got -343

OpenStudy (whpalmer4):

just match things up

OpenStudy (anonymous):

what do you mean?

OpenStudy (whpalmer4):

you can't tell which number corresponds to x and which corresponds to a here? \[7^{-3}\]\[x^{-a}\]

OpenStudy (anonymous):

oh well, x=7 and a= -3?

OpenStudy (whpalmer4):

no, look more closely

OpenStudy (anonymous):

uh

OpenStudy (whpalmer4):

here, I'll make it bigger :-) \[\large 7^{-3}\]\[\large x^{-a}\]

OpenStudy (anonymous):

i dont get it

OpenStudy (whpalmer4):

okay. you have 3 symbols: 7, -, and 3 you have to match them with 3 other symbols: x, -, and a